OpenStudy (anonymous):

A regular polyhedron has 30 edges and 12 vertices. How many faces does it have?

hero (hero):

The general formula to find the face of a polyhedron given the number of edges and vertices is as follows: V - E + F = 2 V = Vertex E = Edges F = Faces In this case, E = 30 and V = 12, therefore we insert those values into the formula and solve for F as follows: 12 - 30 + F = 2 12 + F = 32 F = 32 - 12 F = 20 Therefore, the polyhedron has 20 faces.

OpenStudy (anonymous):

Thanks hero. What if I was only given 1 of the variables, is there a different formula for that?

hero (hero):

That's a good question. If you only have one variable, good luck

OpenStudy (anonymous):

Alright